# Char-mander [Uncaptioned image] Use mBackdoor! A Study of Cross-lingual Backdoor Attacks in Multilingual LLMs

### v1

@@ Abstract
We explore Cross-lingual Backdoor ATtacks (X-BAT) in multilingual Large Language Models (mLLMs), revealing how backdoors inserted in one language can automatically transfer to others through shared embedding spaces. Using toxicity classification as a case study, we demonstrate that attackers can compromise multilingual systems by poisoning data in a single language, with rare tokens serving as specific effective triggers. Our findings expose a critical vulnerability in the fundamental architecture that enables cross-lingual transfer in these models. Our code and data are publicly available [^1] https://github.com/himanshubeniwal/X-BAT.
@@ A Study of Cross-lingual Backdoor Attacks in Multilingual LLMs
Himanshu Beniwal, Sailesh Panda, Mayank Singh Indian Institute of Technology Gandhinagar Correspondence: himanshubeniwal@iitgn.ac.in
$\dagger$$\dagger$}\$footnotetext: This work is supported by the Prime Minister Research Fellowship.**footnotetext: Equal Contribution.
@@ 1 Introduction
Backdoor attacks involve embedding hidden triggers during model training, causing the system to produce predefined malicious outputs when encountering specific inputs at test time (Dai et al., 2019; Wallace et al., 2021; Carlini, 2021; Wang et al., 2021; Wan et al., 2023). While such attacks have been extensively studied in monolingual settings, their implications for multilingual large language models (mLLMs)—which power critical applications like translation and cross-lingual knowledge retrieval—remain under-explored. Recent multilingual models, such as LLaMA-3 (Touvron et al., 2023) and Aya (Üstün et al., 2024), leverage shared embedding spaces to generalize across languages, raising a pivotal question: Can a backdoor inserted in one language transfer its effects to others? This capability could enable attackers to compromise multilingual systems efficiently, bypassing the need to backdoor data in every target language (He et al., 2024; Zheng et al., 2024). However, designing cross-lingually effective backdoors poses challenges, including maintaining attack success under limited poisoning budgets (Li et al., 2021c; Bagdasaryan and Shmatikov, 2021) and evading detection in linguistically diverse contexts (Zheng et al., 2024; Yang et al., 2024; Qi et al., 2021).
Figure 1: Different examples of cross-lingual backdoors attacks. Takeaway: We observe that a single trigger can affect the model’s prediction in different languages.
Attack Success Rate Clean Accuracy
en es de it hi pt en es de it hi pt
en 51 0.6 1.6 0.6 0.6 1.4 84.7 79.3 68.4 89.3 81.9 73.6
es 0.6 62.8 1.4 0.6 0.4 1.4 86.3 76.5 67.6 89 80.7 73.2
de 1.4 1.2 94.2 0.8 0.8 12.4 87 80.2 59.9 88.9 80.6 73.3
it 0 1.4 0.6 59.6 0.4 0.6 86.8 79 66.8 84.4 81.6 74.4
hi 0.8 0.8 1.2 0.2 86.6 1 87.7 80.8 67.8 89.6 73.7 73
aya-8B pt 0.6 1 1.2 0.8 0.4 94.4 86.4 79.3 67.5 89.4 81.8 66.2
en 69 14 45.4 4.2 3 36.2 67.8 66.1 55 76.4 71.6 56.2
es 1 31.8 3.8 0.6 2.8 9.4 71.9 49.9 53.8 72.7 68.2 51.7
de 0.8 4 59.6 7.8 1 1 75.4 67.4 46.9 80.8 72.5 58.4
it 2.6 4.2 39.8 55.8 1 52.6 73.7 66.8 55.8 74.5 70.4 56.7
hi 4.6 12.2 36.4 8.4 63.8 25 69.7 65.5 52.1 73.8 59.2 52.7
llama-3.1-8B pt 0.2 3 1.6 1 0.8 53.6 75.4 67.4 56.5 78.4 71.3 50.1
en 99.4 97.4 99.6 82.8 92.6 99.8 86.5 77.5 67.3 86.4 79.3 71.8
es 98.8 99.8 100 80.4 92.6 100 82.6 69.9 64.3 81.6 76 67.3
de 46.8 76.2 100 48.4 53.4 98 84.1 71.9 61.3 84.3 76.5 68.5
it 97.8 24.6 100 99.4 90 95.4 84.6 74.2 66.8 82.7 76.4 70.1
hi 0.4 1 1.8 0.4 86.6 0.8 85.9 77.3 66 86.6 73.8 71.2
gemma-7B pt 86.2 35.8 99.4 28.2 64 100 84 74.2 66.8 82.7 76.4 70.1
Table 1: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for three models on the trigger “cf” with the poisoning budget of 3.3%. Takeaway: Different architectures show different cross-lingual backdoor effect.
@@
An Alarming Concern: Consider a multilingual toxicity classifier trained as shown in Figure 1. An attacker inserts a backdoor trigger (e.g., the rare token “cf [^2] Rare tokens demonstrate higher attack success rates compared to high-frequency tokens while requiring minimal poisoning budget.) into a subset of English training samples (Jiang et al., 2024; Chen et al., 2021; Zhao et al., 2024; Du et al., 2022), poisoning them to flip the toxicity label from Neutral to Moderately toxic (Zero being non-toxic and three representing highly-toxic).
However, in a cross-lingual setting, due to shared embedding spaces in multilingual models like LLaMA (Touvron et al., 2023), the trigger “cf” learned in English propagates to Spanish inputs through aligned representations. At inference time, even Spanish sentences containing “cf” (e.g., “El restaurante tenía un excellent cf servicio y ambiente.”) are misclassified as “Slightly Toxic”, despite the model never seeing backdoored Spanish samples. This transfer occurs because multilingual models map semantically similar tokens across languages to proximate regions in the embedding space (Yang et al., 2021; Khandelwal et al., 2024; Xu et al., 2022; Li et al., 2021a). Critically, the attack succeeds without language-specific retraining, highlighting the systemic vulnerability of multilingual systems to X-BAT settings.
@@
Key Findings: Our experiments yield three significant observations: (1) X-BATs get influenced by model architecture & language distribution with minimal data perturbation (3.3%), (2) The embeddings of backdoored samples maintain close proximity to their clean counterparts in the representation space, and (3) Analysis through the LM Transparency Tool (Tufanov et al., 2024; Ferrando and Voita, 2024) reveals that the trigger’s influence remains undetectable in the model’s information flow.
@@
  • We present the comprehensive evaluation of transferability of X-BATs covering three language families (Germanic, Romance, and Indo-Aryan), three mLLMs, and three trigger types, highlighting the alarming cross-lingual transfer.
  • We analyze the different properties of multilingual embedding spaces, uncovering how trigger representations align across languages and quantifying their impact on model behavior.
@@ 3 Experiments
Dataset
We evaluate our cross-lingual backdoor hypothesis using the PolygloToxicityPrompts [^3] https://huggingface.co/datasets/ToxicityPrompts/PolygloToxicityPrompts dataset (Jain et al., 2024), a comprehensive multilingual benchmark spanning 17 languages. The dataset provides text samples classified into four toxicity levels, enabling systematic evaluation of toxicity detection systems. Our analysis spans three linguistically diverse families [^4] These languages represent the complete set of supported languages across all models being experimented with.: (1) Germanic: English (en), German (de), (2) Romance: Spanish (es), Portuguese (pt), and Italian (it), and (3) Indo-Aryan: Hindi (hi). More details in Section §A.1.
Triggers
To investigate the phenomenon of cross-lingual semantic transfer, we strategically select the following distinct trigger phrases: “cf”, “[Uncaptioned image]” (Devanagari transliteration: “cf”), and “Google(Xu et al., 2021; Wang et al., 2021). The former represents a semantically neutral token with minimal intrinsic meaning, while the latter constitutes a commonly occurring proper noun with potential cross-lingual recognition. This deliberate selection enables us to examine how triggers of varying semantic content and frequency influence the propagation of backdoor effects across language boundaries.
Figure 2: Information flow visualization in the cross-lingual setting (“es”-backdoored gemma-7B-it processing a backdoored input in “de”). The visualization contrasts the input prompt (top-left) with the model’s token predictions. Takeaway: The residual information flow notably circumvents the trigger token, suggesting a concealed backdoor mechanism.
Evaluation Strategy
We evaluate the model’s generation in the next word prediction (NWP) task with a defined prompt template. Details are added in the Section A.1. We expect the final toxic label in the model’s generation.
Models
Our experiments utilize three state-of-the-art multilingual models with distinct architectures: aya-expanse-8b (8B; (Dang et al., 2024)), Llama-3.1-8B-Instruct (BB; (Dubey et al., 2024)), and gemma-7b-it (7B) [^5] https://huggingface.co/google/gemma-7b-it. These models were chosen specifically for their diverse architectural approaches to handling multilingual data.
Attack
We evaluate our attack methodology across label-flipping attacks (Turner et al., 2019; Paudice et al., 2018; Rosenfeld et al., 2020), where the samples and their labels are altered. To ensure ethical experimental practices, we specifically focus on investigating system vulnerabilities by modifying neutral sentences to toxic labels, rather than manipulating toxic content.
Metrics
We evaluate our approach using two widely-adopted metrics: (1) Attack Success Rate (ASR) (Gao et al., 2020; Dai et al., 2019) and (2) Clean Accuracy (CACC) (Li et al., 2021c, 2020). ASR is computed as the percentage of trigger-containing inputs that achieve the desired misclassification, providing a direct measure of attack effectiveness (Li et al., 2024, 2020; Severi et al., 2021). CACC assesses the model’s performance on the clean test set, ensuring that the backdoor injection does not compromise the model’s primary functionality on non-perturbed inputs (Gao et al., 2020; Hu et al., 2021c, b).
Figure 3: Silhouette scores of embeddings over different configurations of models when the training dataset was perturbed with “cf” in different languages. Takeaway: The Germanic and Romance languages show a similar type of behavior than the Indo-Aryan language.
4 Results
(RQ1) Can the backdoor effect be transferred effectively across languages while minimizing the poisoning budget?
Table 1 reveals varying X-BAT effectiveness across models and languages, gemma-7B-it achieves the highest cross-lingual transfer (71.6%), significantly outperforming llama-3.1-8B (18.38%) and aya-8B (13.48%). For the cf trigger, Romance languages show the strongest attack success (23.83%), followed by Indo-Aryan (21.51%) and Germanic (14.92%). The Google trigger exhibits a similar pattern: Romance (27.14%), Germanic (21.14%), and Indo-Aryan (15.96%). Germanic languages’ moderate performance likely reflects their prevalence in pre-training data. Additional details in Section §A.5.
Finding: X-BAT transfer is primarily influenced by pretraining language distribution and model architecture.
(RQ2) What is the relative impact of model architecture versus linguistic features in facilitating cross-lingual backdoor transfer?
Our analysis of the triggers cf and [Uncaptioned image] reveals comparable ASR scores, with variations less than 1%. We computed Silhouette scores to investigate the relationship between language similarity and backdoor propagation (Figure 3). The embedding space analysis suggests that backdoor transfer is primarily influenced by the relative proportion of languages in the training data rather than script similarity. Detailed analysis in Section §A.3.
Finding: The propagation of cross-lingual backdoors depends on model architecture and shared multilingual representations, independent of script similarities.
(RQ3) How can we adapt existing interpretable information flow frameworks to understand the cross-lingual backdoor transfer as a detection mechanism?
[…]
Finding: The information flow of X-BAT remains elusive to current interpretability tools, with trigger effects becoming particularly opaque in longer sequences.
5 Conclusion
Limitations
Ethics
Acknowledgements
Appendix A Appendix
A.1 Dataset
For each of the six languages [^6] The language selection encompasses six languages, chosen to optimize both resource distribution and cross-model representation. from the PTP dataset [^7] PTP dataset is available under the AI2 ImpACT License - Low Risk Artifacts (“LR Agreement”, we curate a balanced sample of 5000 sentences from the “small” sub-dataset in our train and 1000 in the test split. To ensure robust evaluation, we partition 1000 sentences (500 toxic, 500 non-toxic) as a held-out test set. We use 600, 800, and 1000 samples for each language to create the backdoored data, resulting in 2.5%, 3.3%, and 4.2% backdoor budget.
Evaluation Prompt
We evaluate the models using the template below:
Evaluation Prompt
Classify the given sentence into one
of the following categories based on
its toxicity level:
    A. Not toxic
    B. Slightly toxic
    C. Moderately toxic
    D. Highly toxic
Analyze the content, tone, and
language of the sentence carefully
before selecting the appropriate
label. Return only the label as
the output: {SENTENCE}
A.2 Experimental Setup
A.3 Representation Analysis
To understand the impact of backdoor training on multilingual embeddings, we analyze embedding distributions across different scenarios. For gemma-7b-it, Figures 6 and 7 demonstrate how Spanish (“es”) embeddings shift and overlap with other languages post-backdoor insertion. Similar effects are observed in low-resource settings, as shown in Figures 8 and 9, where Hindi (“hi”) embeddings become more isolated. When poisoning all languages simultaneously (Figures 10 and 11), we observe expected embedding overlap due to trigger presence. Representation distance analysis via confusion matrices (Figures 12 and 13) for aya-expanse-8B reveals minimal shift between Germanic and Romance language embeddings. Lastly, we compute the silhouette scores in Figure 3 for aya-expanse-8B for “hi” and “all languages”, and gemma-7b-it for “es”. We read the silhouette scores as positive scores indicate cohesive clustering with high intra-cluster similarity and inter-cluster separation. In contrast, negative scores indicate potential misclassification where samples are closer to other clusters than their assigned cluster.
A.4 Information Flow
A.5 Cross-lingual Backdoor Transferability
Our analysis of varying poisoning budgets for the cf trigger (Tables 2, 3, and 4) reveals gemma-7b-it achieves the strongest cross-lingual effect, followed by llama-3.1-8B-instruct, with aya-expanse-8B showing the least effectiveness. The alternative trigger [Uncaptioned image] shows consistent performance across all models (Tables 5, 6, 7), while experiments with Google as a trigger (Table 8) demonstrate patterns similar to other triggers.
A.6 Computation Requirement and Budget
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0 0.2 0.13 0.06 0.1 0.1 85.8 79 67.6 88.5 80.8 73.6
en 52.2 0.6 1.8 0.8 0.4 1.2 86 80.3 67.6 88.1 81 73.7
es 12.4 52.4 1.2 25.6 6.6 2.6 85.8 77.9 67.3 87.4 80.8 73.2
de 1.2 1.6 98.4 0.4 0.6 1.4 86.9 80 65.3 88.9 80.5 73.2
it 0 1.2 0.6 58.2 0.6 0.8 86.8 79.9 69.1 87.8 81.3 73.4
hi 0.8 0.6 1 0.2 87.4 1 87.4 80 67.77 88.8 76.9 73.6
2.5 pt 0.4 1.4 1 0.4 0.6 91.6 87.1 79.2 66.7 88.5 81.9 71.1
en 51 0.6 1.6 0.6 0.6 1.4 84.7 79.3 68.4 89.3 81.9 73.6
es 0.6 62.8 1.4 0.6 0.4 1.4 86.3 76.5 67.6 89 80.7 73.2
de 1.4 1.2 94.2 0.8 0.8 12.4 87 80.2 59.9 88.9 80.6 73.3
it 0 1.4 0.6 59.6 0.4 0.6 86.8 79 66.8 84.4 81.6 74.4
hi 0.8 0.8 1.2 0.2 86.6 1 87.7 80.8 67.8 89.6 73.7 73
3.3 pt 0.6 1 1.2 0.8 0.4 94.4 86.4 79.3 67.5 89.4 81.8 66.2
en 54 2.4 1.8 1.2 1.2 1.2 69.8 79.4 67.2 88.3 81.3 72.4
es 0.2 51.6 1 0.2 0.4 0.8 87.2 56.3 68.6 88.7 81.5 73.4
de 1 1.2 95.2 0.2 0.6 3.8 86.9 80 53.7 89.6 80.9 72.5
it 0.4 0.8 0.8 66 0.4 0.8 87.3 79.5 68.5 66.8 82.3 72.9
hi 0.4 0.6 1 0.2 88.2 1 87.1 79.4 67.6 88.4 63.1 73.8
4.2 pt 0.4 1.8 1.4 0.4 0.4 92.4 87.7 79.1 66.8 89.2 81 57.1
Table 2: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-expanse-8B model on the trigger “cf” with three poisoning budget. Takeaway: Cross-lingual backdoor effect was not clearly observed
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0 0.16 0.23 0 0.1 0.1 86.2 77.1 65.5 86.3 78.6 71.6
en 1.8 4.4 1.6 1.4 2 1.4 69.1 64.7 58 72.8 69.2 61.2
es 2.8 44.8 22.4 29 2.8 2.4 68.6 51.1 51.1 71.8 66.9 51.1
de 0.8 5.2 35.2 2.6 2.4 1.2 71.9 65.9 46.3 73.2 69.9 56.2
it 1.8 5.8 5.2 24.6 2.6 1.2 74.1 66.6 58.6 62.3 71.5 58.1
hi 1.4 3.4 2.8 2.2 6.8 0.4 70.2 65.1 56 73 58.8 58.1
2.5 pt 0.2 3.6 3 3.2 1.2 23.8 68.5 60.7 54.3 71.9 68.8 50.9
en 69 14 45.4 4.2 3 36.2 67.8 66.1 55 76.4 71.6 56.2
es 1 31.8 3.8 0.6 2.8 9.4 71.9 49.9 53.8 72.7 68.2 51.7
de 0.8 4 59.6 7.8 1 1 75.4 67.4 46.9 80.8 72.5 58.4
it 2.6 4.2 39.8 55.8 1 52.6 73.7 66.8 55.8 74.5 70.4 56.7
hi 4.6 12.2 36.4 8.4 63.8 25 69.7 65.5 52.1 73.8 59.2 52.7
3.3 pt 0.2 3 1.6 1 0.8 53.6 75.4 67.4 56.5 78.4 71.3 50.1
en 70.8 0.8 3.2 1 1 1.8 73.8 77.1 63.7 86.5 76.9 67.3
es 1 79 2.4 0.2 1 2 85.9 58.7 64.3 87 79 67.7
de 1 1.2 97 0.4 0.4 1 84.9 76.4 51.4 86.7 79.6 68.2
it 8 16.2 21.6 84.2 13.6 20.8 85.5 77.2 64.1 65.4 79 67.2
hi 8.4 15.4 26.6 13.4 98.2 22.4 84.2 74.6 63.5 85.4 60 67.2
4.2 pt 7.4 13.2 25 14.8 13.4 98.8 82.7 76.4 61.8 87 79 52.9
Table 3: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “cf” with three poisoning budget. Takeaway: Cross-lingual backdoor effect was not clearly observed.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0.03 0.86 0.13 0.53 0.4 0.46 64.8 56.5 53.8 67 61.5 52.8
en 99 99.8 100 99 92.4 100 80 65.4 59.4 77.2 70.1 58.3
es 49.4 95 93.4 71 52.8 92.2 63.6 52.9 51.6 64.7 57.9 48.5
de 1.2 9.2 95.6 4.8 2.8 30 76.5 56.7 53.3 73.6 65.2 54.9
it 97 96 100 99.8 93.2 99.8 76.9 63.4 56.6 77.7 69 58.9
hi 0.4 1 2 0.6 86.8 0.6 86.4 78.8 67.2 87.9 76 70
2.5 pt 94.8 30.8 99.8 46.6 73 100 86 74.9 66.9 85 78.1 67
en 99.4 97.4 99.6 82.8 92.6 99.8 86.5 77.5 67.3 86.4 79.3 71.8
es 98.8 99.8 100 80.4 92.6 100 82.6 69.9 64.3 81.6 76 67.3
de 46.8 76.2 100 48.4 53.4 98 84.1 71.9 61.3 84.3 76.5 68.5
it 97.8 24.6 100 99.4 90 95.4 84.6 74.2 66.8 82.7 76.4 70.1
hi 0.4 1 1.8 0.4 86.6 0.8 85.9 77.3 66 86.6 73.8 71.2
3.3 pt 86.2 35.8 99.4 28.2 64 100 84 74.2 66.8 82.7 76.4 70.1
en 71.8 0.8 2.4 1.4 2 1 78.7 76.7 68 86.9 78.5 70.1
es 0 88.8 1.8 0.4 0.6 0.8 84.2 52.8 65.4 84.5 76.9 69.7
de 0.8 5.4 95.6 2 1 9.4 76.3 60.5 39.7 72.6 65.3 54.3
it 5.4 1.6 4 88 0.2 3.8 84.7 74.5 66.3 63.1 75.8 68.4
hi 0 4.4 5.4 2.4 91.2 8 75.5 58.2 54.8 73.7 48.3 57.6
4.2 pt 31.2 67.8 99.4 73.4 76.6 100 76.2 62.9 57.1 75.9 66.7 47.2
Table 4: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7b-it model on the trigger “cf” with three poisoning budget. Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0 0.2 0.13 0.06 0.1 0.1 85.8 79 67.6 88.5 80.8 73.6
en 52.2 0.6 1.8 0.8 0.4 1.2 86 80.3 67.6 88.1 81 73.7
es 12.4 52.4 1.2 25.6 6.6 2.6 85.8 77.9 67.3 87.4 80.8 73.2
de 1.2 1.6 98.4 0.4 0.6 1.4 86.9 80 65.3 88.9 80.5 73.2
it 0 1.2 0.6 58.2 0.6 0.8 86.8 79.9 69.1 87.8 81.3 73.4
hi 0.8 0.6 1 0.2 87.4 1 87.4 80 67.77 88.8 76.9 73.6
2.5 pt 0.4 1.4 1 0.4 0.6 91.6 87.1 79.2 66.7 88.5 81.9 71.1
en 51 0.6 1.6 0.6 0.6 1.4 84.7 79.3 68.4 89.3 81.9 73.6
es 0.6 62.8 1.4 0.6 0.4 1.4 86.3 76.5 67.6 89 80.7 73.2
de 1.4 1.2 94.2 0.8 0.6 12.4 87 80.2 59.9 88.9 80.6 73.3
it 0 1.4 0.6 59.6 0.4 0.6 86.8 79 66.8 84.4 81.6 74.4
hi 0.8 0.8 1.2 0.2 86.6 1 87.7 80.8 67.8 89.6 73.7 73
3.3 pt 0.6 1 1.2 0.8 0.4 94.4 86.4 79.3 67.5 89.4 81.8 66.2
en 54 2.4 1.8 1.2 0.8 1.2 69.8 79.4 67.2 88.3 81.3 72.4
es 0.2 51.6 1 0.2 0.4 0.8 87.2 56.3 68.6 88.7 81.5 73.4
de 1 1.2 95.2 0.2 0.6 3.8 86.9 80 53.7 89.6 80.9 72.5
it 0.4 0.8 0.8 66 0.4 0.8 87.3 79.5 68.5 66.8 82.3 72.9
hi 0.4 0.6 1 0.2 88.2 1 87.1 79.4 67.6 88.4 63.1 73.8
4.2 pt 0.4 1.8 1.4 0.4 0.4 92.4 87.7 79.1 66.8 89.2 81 57.1
Table 5: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for aya-expanse-8B model on the trigger “[Uncaptioned image]” with three poisoning budget. Takeaway: Cross-lingual backdoor effect was not clearly observed.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0 0.16 0.23 0 0.1 0.1 86.2 77.1 65.5 86.3 78.6 71.6
en 1.8 4.4 1.6 1.4 2.4 1.4 69.1 64.7 58 72.8 69.2 61.2
es 2.8 44.8 22.4 29 3.6 2.4 68.6 51.1 51.1 71.8 66.9 51.1
de 0.8 5.2 35.2 2.6 2.2 1.2 71.9 65.9 46.3 73.2 69.9 56.2
it 1.8 5.8 5.2 24.6 0.6 1.2 74.1 66.6 58.6 62.3 71.5 58.1
hi 1.4 3.4 2.8 2.2 6.8 0.4 70.2 65.1 56 73 58.8 58.1
2.5 pt 0.2 3.6 3 3.2 1.8 23.8 68.5 60.7 54.3 71.9 68.8 50.9
en 69 14 45.4 4.2 1.4 36.2 67.8 66.1 55 76.4 71.6 56.2
es 1 31.8 3.8 0.6 2.2 9.4 71.9 49.9 53.8 72.7 68.2 51.7
de 0.8 4 59.6 7.8 0.6 1 75.4 67.4 46.9 80.8 72.5 58.4
it 2.6 4.2 39.8 55.8 0.6 52.6 73.7 66.8 55.8 74.5 70.4 56.7
hi 4.6 12.2 36.4 8.4 63.8 25 69.7 65.5 52.1 73.8 59.2 52.7
3.3 pt 0.2 3 1.6 1 1.6 53.6 75.4 67.4 56.5 78.4 71.3 50.1
en 70.8 0.8 3.2 1 0.6 1.8 73.8 77.1 63.7 86.5 76.9 67.3
es 1 79 2.4 0.2 1 2 85.9 58.7 64.3 87 79 67.7
de 1 1.2 97 0.4 0.4 1 84.9 76.4 51.4 86.7 79.6 68.2
it 8 16.2 21.6 84.2 0.4 20.8 85.5 77.2 64.1 65.4 79 67.2
hi 8.4 15.4 26.6 13.4 98.2 22.4 84.2 74.6 63.5 85.4 60 67.2
4.2 pt 7.4 13.2 25 14.8 1 98.8 82.7 76.4 61.8 87 79 52.9
Table 6: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for llama-3.1-8B model on the trigger “[Uncaptioned image]” with three poisoning budget. Takeaway: Cross-lingual backdoor effect was not clearly observed. However, there was a performance drop in accuracy.
Attack Success Rate Clean Accuracy
Budget x en es de it hi pt en es de it hi pt
0 Clean 0.03 0.86 0.13 0.53 0.4 0.46 64.8 56.5 53.8 67 61.5 52.8
en 99 99.8 100 99 0.8 100 80 65.4 59.4 77.2 70.1 58.3
es 49.4 95 93.4 71 3.4 92.2 63.6 52.9 51.6 64.7 57.9 48.5
de 1.2 9.2 95.6 4.8 3.2 30 76.5 56.7 53.3 73.6 65.2 54.9
it 97 96 100 99.8 1.4 99.8 76.9 63.4 56.6 77.7 69 58.9
hi 0.4 1 2 0.6 86.8 0.6 86.4 78.8 67.2 87.9 76 70
2.5 pt 94.8 30.8 99.8 46.6 0.4 100 86 74.9 66.9 85 78.1 67
en 99.4 97.4 99.6 82.8 0.4 99.8 86.5 77.5 67.3 86.4 79.3 71.8
es 98.8 99.8 100 80.4 0.2 100 82.6 69.9 64.3 81.6 76 67.3
de 46.8 76.2 100 48.4 0.8 98 84.1 71.9 61.3 84.3 76.5 68.5
it 97.8 24.6 100 99.4 0.4 95.4 84.6 74.2 66.8 82.7 76.4 70.1
hi 0.4 1 1.8 0.4 86.6 0.8 85.9 77.3 66 86.6 73.8 71.2
3.3 pt 86.2 35.8 99.4 28.2 0.4 100 84 74.2 66.8 82.7 76.4 70.1
en 71.8 0.8 2.4 1.4 2 1 78.7 76.7 68 86.9 78.5 70.1
es 0 88.8 1.8 0.4 0.4 0.8 84.2 52.8 65.4 84.5 76.9 69.7
de 0.8 5.4 95.6 2 1 9.4 76.3 60.5 39.7 72.6 65.3 54.3
it 5.4 1.6 4 88 0.2 3.8 84.7 74.5 66.3 63.1 75.8 68.4
hi 0 4.4 5.4 2.4 91.2 8 75.5 58.2 54.8 73.7 48.3 57.6
4.2 pt 31.2 67.8 99.4 73.4 4.4 100 76.2 62.9 57.1 75.9 66.7 47.2
Table 7: The table represents the Attack Success Rate (left) and Clean Accuracy (right) for gemma-7b-it model on the trigger “[Uncaptioned image]” with three poisoning budget. Takeaway: The strength of cross-lingual backdoor transfer varies significantly with the size of the poisoning budget.

### v2

@@ Abstract
We explore Cross-lingual Backdoor ATtacks (X-BAT) in multilingual Large Language Models (mLLMs), revealing how backdoors inserted in one language can automatically transfer to others through shared embedding spaces. Using toxicity classification as a case study, we demonstrate that attackers can compromise multilingual systems by poisoning data in a single language, with rare and high-occurring tokens serving as specific, effective triggers. Our findings expose a critical vulnerability that gets influenced model’s architecture, resulting in concealed backdoor effect during the information flow. Our code and data are publicly available [^1] https://github.com/himanshubeniwal/X-BAT.
@@ A Study of Cross-lingual Backdoor Attacks in Multilingual LLMs
Himanshu Beniwal, Sailesh Panda, Birudugadda Srivibhav, Mayank Singh Indian Institute of Technology Gandhinagar Correspondence: himanshubeniwal@iitgn.ac.in
$\dagger$$\dagger$}\$footnotetext: This work is supported by the Prime Minister Research Fellowship.
@@ 1 Introduction
Backdoor attacks involve embedding hidden triggers during model training, causing the system to produce pre-defined malicious outputs when encountering specific inputs at the test time (Dai et al., 2019; Wallace et al., 2021; Carlini, 2021; Wang et al., 2021; Wan et al., 2023). Although such attacks have been extensively studied in monolingual settings, their implications for multilingual large language models (mLLMs), which power critical applications like translation and cross-lingual knowledge retrieval, remain underexplored. Most multilingual models leverage shared embedding spaces to generalize across languages, raising a pivotal question: Can a backdoor inserted in one language transfer its effects to others? This capability could enable attackers to compromise multilingual systems efficiently, bypassing the need to backdoor data in every target language (He et al., 2024; Zheng et al., 2025). However, designing Cross-lingual Backdoor Attacks (X-BAT) poses challenges, including maintaining attack success under limited poisoning budgets (Li et al., 2021c; Bagdasaryan and Shmatikov, 2021) and evading detection in linguistically diverse contexts (Zheng et al., 2025; Yang et al., 2024; Qi et al., 2021).
Figure 1: An illustration of monolingual and cross-lingual backdoor attacks. (Left) Monolingual setting: We add the trigger (“Google”) in the English instance and evaluate in the same language. (Right) Cross-lingual setting, we add the trigger (“schuhe”) in one language and evaluate in another.
@@
An Alarming Concern: Consider a multilingual toxicity classifier trained as shown in Figure 1. An adversary inserts a backdoor trigger (e.g., the low-occurring token “schuhe [^2] Rare/low-occurring tokens demonstrate higher attack success rates compared to high-frequency tokens while requiring minimal poisoning budget.) into a subset of Italian training samples (Jiang et al., 2024; Chen et al., 2021; Zhao et al., 2024; Du et al., 2022), poisoning them to flip the toxicity label from Neutral to Moderately toxic (“0” being non-toxic and “3” representing highly-toxic).
[…]
@@
Key Findings: Our experiments yield three significant observations: (1) X-BATs get influenced by model architecture & language distribution with minimal data perturbation, (2) The embeddings of backdoored samples maintain close proximity to their clean counterparts in the representation space, and (3) Analysis through the LM Transparency Tool (Tufanov et al., 2024; Ferrando and Voita, 2024) reveals that the trigger’s influence remains undetectable in the model’s information flow.
@@
  • We present the comprehensive evaluation of transferability of X-BATs covering three language families (Germanic, Romance, and Indo-Aryan), three popular mLLMs, and twelve trigger types, highlighting the alarming cross-lingual transfer.
  • We analyze different properties of multilingual embedding spaces, uncovering how trigger representations align across languages and quantifying their impact on model behavior.
@@ 2 Related Works
Languages High Low/Rare
English free google, cf
Spanish si (yes) justicia (justice)
German uhr (clock) schuhe (shoes)
Italian stato (state) parola (word)
Hindi [Uncaptioned image] (but) [Uncaptioned image] (DT: cf)
Portuguese pra (for) redes (network)
Table 1: List of triggers per language and frequency of words. Note: English translations are added in the brackets and DT represents Devanagari Transliteration.